# 00footnotetext: *Equal contribution

### v1

@@ 1 Abstract
In this report, we present ML-Dev-Bench [^1] https://github.com/ml-dev-bench/ml-dev-bench, a benchmark aimed at testing agentic capabilities on applied Machine Learning development tasks. While existing benchmarks focus on isolated coding tasks or Kaggle-style competitions, ML-Dev-Bench tests agents’ ability to handle the full complexity of ML development workflows. The benchmark assesses performance across critical aspects including dataset handling, model training, improving existing models, debugging, and API integration with popular ML tools. We evaluate three agents - ReAct, Openhands, and AIDE - on a diverse set of 25 tasks, providing insights into their strengths and limitations in handling practical ML development challenges.
@@ 2 Introduction
Benchmarks like ML-Bench [6], test agents’ abilities to generate code and commands to interact with popular ML repositories, while MLE-Bench [2] and MLAgentBench [4] focus on Kaggle-style tasks to evaluate the iterative and open-ended nature of ML development. However, real-world ML development extends far beyond that, including the complexity of working on top of existing codebases and models, integrating with third-party tools, debugging complex issues that span multiple components of the ML pipeline and understanding and balancing trade-offs like model performance and cost to come up with optimal design.
ML-Dev-Bench addresses this gap by providing a comprehensive evaluation framework that tests an agent’s ability to handle real-world ML development scenarios. Our benchmark is particularly relevant as ML development increasingly relies on large language models and AI agents to assist developers. Understanding the capabilities and limitations of these agents in handling practical ML development tasks is crucial for their effective deployment in production environments.
@@ 3 Benchmark Design
ML-Dev-Bench comprises of 25 carefully designed tasks that evaluate various aspects of ML development. These tasks are structured to assess both specific technical capabilities (like handling datasets, model implementation) and broader problem-solving skills (like model training and performance improvement) that are essential in real-world ML development. The tasks span several key categories of ML development shown in Table 1:
@@
    Performance optimization challenges agents to improve baseline implementations through iterative experimentation and hypothesis testing.
Category ReAct-Sonnet OH-Sonnet Aide-4o ReAct-4o
Dataset Handling 100% (3/3) 100% (3/3) 33% (1/3) 0% (0/3)
Model Training 67% (4/6) 83% (5/6) 33% (2/6) 50% (3/6)
Debugging 67% (4/6) 67% (4/6) 33% (2/6) 16% (1/6)
API Integration 100% (1/1) 100% (1/1) 0% (0/1) 100% (1/1)
Model Implementation 50% (2/4) 50% (2/4) 0% (0/4) 0% (0/4)
Performance 0% (0/5) 0% (0/5) 0% (0/5) 0% (0/5)
Overall 56% (14/25) 60% (15/25) 20% (5/25) 20% (5/25)
Table 2: Category-wise Success Rates Across AI Agents
@@ 4.1 Evaluation Task
Each evaluation task consists of a task description, a set of input code and data files, and a validation logic which checks agent generated outputs and artifacts for correctness. Depending on the type of task, we implement various types of validation checks including
@@ 5 Agent Setup
    Openhands: Openhands [7] is a popular open-source coding agent with state-of-the-art performance on SWE-Bench-Full [5]. We used Openhands agent v0.21.1 and customized the runtime build to install common ML frameworks listed above. We tested the agent with Claude Sonnet 3.5 10-2022 model which is the current best performing model with the agent on SWE Bench. The agent was allowed to run for a maximum of 50 steps.
@@
Task Category ReAct-Sonnet OH-Sonnet Aide-4o ReAct-4o
Dataset download - Noisy Imagenette Dataset Handling × ×
Dataset download - dataset does not exist Dataset Handling ×
Dataset preprocessing Dataset Handling × ×
Pretrained model download - Torchvision Model Training
Pretrained model download - HuggingFace Model Training × ×
Vision finetuning - classification Model Training × ×
Overfit on small dataset Model Training
Large training logs Model Training × ×
CIFAR10 Training Model Training × × × ×
Fix problems in model and dataloader Debugging × × × ×
Model forward pass - shape mismatches Debugging
Model Training - shape mismatches Debugging ×
NaN losses Debugging × ×
Correct norm for pretrained model Debugging × ×
TinyBERT Eval Debugging × × × ×
Wandb integration API Integration ×
ChannelViT - Easy Model Implementation × ×
ChannelViT Model Implementation × ×
ChannelViT - No tests Model Implementation × × × ×
VAR implementation Model Implementation × × × ×
Improve CIFAR-10 baseline - existing model ckpt Performance × × × ×
Noisy Imagenette Performance × × × ×
CIFAR-10 long tailed Performance × × × ×
Segmentation Performance × × × ×
BoolQ Performance × × × ×
Success Rate 56% (14/25) 60% (15/25) 20% (5/25) 20% (5/25)
Table 3: Performance Comparison Across AI Agents
@@ 6 Performance Comparison
Performance of the agents across different task categories, Table 2 and individual tasks Table 3 reveals a consistent pattern. Performance decreases as tasks become more open-ended and complex. The success rates are highest in well-defined categories like dataset handling and basic debugging, but drop significantly in performance optimization tasks where no agent succeeded.
OpenHands-Sonnet (OH-Sonnet) and ReAct-Sonnet are the two best performing agents with 60% and 56% success rate respectively, while AIDE-4o and React-4o achieve 20% success rate. Both OH and ReAct Sonnet do well in structured tasks like dataset download, model training setup and debugging tasks with clear instructions but struggle in open-ended performance improvement and long-running training tasks.
Across the 14 common successful tasks between OH-Sonnet and ReAct-Sonnet, we did not observe any discernable trends in cost. In long running tasks OH-Sonnet used more tokens on average, especially in debugging and training tasks, while ReAct had higher usage in model implementation tasks like ChannelViT. Even in tasks with higher token usage, the costs dont scale proportionally due to efficient prompt caching in Openhands.
@@ 6.1 ReAct-Sonnet
ReAct-Sonnet achieved a success rate of 56% (14/25 tasks), demonstrating strong performance in specific, well-defined tasks but struggling with more complex scenarios. The agent excelled in dataset handling (3/3 tasks), basic model training (4/6 tasks), and debugging tasks with clear specifications (4/6 tasks). However, its performance degraded significantly in model implementation and performance optimization categories, where tasks required more open-ended problem-solving.
A notable strength was ReAct-Sonnet’s systematic approach to debugging when provided with specific instructions and test cases. However, the agent showed several limitations listed below:
@@
    Premature task termination in long-running training scenarios, where it doesnt wait until completion of long running tasks like training.
@@
    In long running tasks it fails to successfully implement sub-tasks which it handled correctly in insolation. For example, it correctly handles the noisy imagenette dataset setup in dedicated download tasks but fails the same operation when it’s part of the longer training pipeline.
@@ 6.2 ReAct-4o
ReAct-4o had some success with tasks with well-defined specifications (WandB logging, downloading a specific model from Torchvision), and certain debugging tasks. However it did struggle on other tasks in the same categories. It also failed on the relatively easier tasks like dataset download due to not following instructions, ran into indentation errors while attempting to debug code and failing to produce output artifacts as required by certain tasks.
6.3 OpenHands-Sonnet
OpenHands-Sonnet demonstrated the highest success rate at 60% (15/25 tasks), showing robust performance across most categories. The agent successfully completed all dataset handling tasks (3/3) and showed strong performance in model training (5/6) and debugging (4/6).
The agent particularly excelled in structured tasks and showed better persistence in long-running operations. However, it struggled with performance optimization tasks (0/5), indicating limitations in open-ended problem-solving scenarios requiring iterative improvement.
@@ 6.4 Aide-4o
Aide-4o had a 20% success rate (5/25 tasks), demonstrating limitations across most categories. The agent managed to complete some basic dataset handling and debugging tasks but struggled with model training (2/6) and completely failed in model implementation and performance optimization categories.
@@
Task Token Cost (\$) Total Tokens
ReAct OH ReAct OH
Vision finetuning - classification 0.176 0.124 67,034 54,434
ChannelViT 1.06 0.215 338,055 177,182
ChannelViT - Easy 1.090 0.318 352,141 267,508
ChannelViT - No tests 0.091 0.121 33,208 53,475
Dataset download - dataset does not exist 0.018 0.051 13,735 16,125
Dataset preprocessing 0.078 0.103 27,210 42,277
Model forward pass - shape mismatches 0.069 0.075 27,629 44,396
Pretrained model download - HuggingFace 0.096 0.063 36,592 24,738
CIFAR-10 long tailed 0.089 0.334 29,659 351,943
Fix problems in model and dataloader 0.376 0.556 146,826 785,182
NaN losses 0.124 0.212 40,385 193,801
Dataset download - Noisy Imagenette 0.129 0.068 67,426 25,429
Correct norm for pretrained model 0.380 0.265 136,831 313,149
Overfit on small dataset 0.093 0.133 25,642 52,791
Large training logs 0.023 0.185 35,189 114,903
Segmentation 0.118 0.383 39,662 395,127
Pretrained model download - Torchvision 0.058 0.044 22,289 30,596
CIFAR10 Training 0.209 0.288 71,409 253,445
Model Training - shape mismatches 0.289 0.147 115,568 92,262
Add implementation - VAR 0.051 0.139 12,473 59,863
Wandb integration 0.155 0.258 65,810 266,738
TinyBERT Eval 0.313 0.573 121,773 762,780
BoolQ 0.343 0.650 131,458 993,175
Improve CIFAR-10 baseline - existing model ckpt 0.115 0.401 29,232 366,660
Noisy Imagenette 0.582 0.288 192,275 253,445
Table 4: Comparison of Token Metrics between ReAct-Sonnet and OpenHands-Sonnet
@@ 7 Conclusion
We presented ML-Dev-Bench, a benchmark focused on ML development workflows consisting of 25 tasks. We evaluated 3 agents on this benchmark - ReAct (with Claude Sonnet and GPT-4o), Openhands and AIDE; Openhands with Claude Sonnet performed the best out of these. Future work can involve analysing the impact of scaling compute on these agents; computing variance in success metrics across multiple runs; including reasoning models such DeepSeek-R1, O-1/O-3; and expanding the problem categories to include areas such as label collection. We open-source the evaluation framework for the benefit of the broader community.

### v2

@@ 1 Abstract
In this report, we present ML-Dev-Bench [^1] https://github.com/ml-dev-bench/ml-dev-bench, a benchmark aimed at testing agentic capabilities on applied Machine Learning development tasks. While existing benchmarks focus on isolated coding tasks or Kaggle-style competitions, ML-Dev-Bench tests agents’ ability to handle the full complexity of ML development workflows. The benchmark assesses performance across critical aspects including dataset handling, model training, improving existing models, debugging, and API integration with popular ML tools. We evaluate three agents - ReAct, Openhands, and AIDE - on a diverse set of 30 tasks, providing insights into their strengths and limitations in handling practical ML development challenges. We open source the benchmark for the benefit of the community.
@@ 2 Introduction
Benchmarks such as ML-Bench [6], test agents’ abilities to generate code and commands to interact with popular ML repositories, while MLE-Bench [2] and MLAgentBench [4] focus on Kaggle-style tasks to evaluate the iterative and open-ended nature of ML development. However, real-world ML development extends far beyond that, including the complexity of working on top of existing codebases and models, integrating with third-party tools, debugging complex issues that span multiple components of the ML pipeline, and understanding and balancing trade-offs like model performance and cost to come up with optimal design.
ML-Dev-Bench addresses this gap by providing a comprehensive evaluation framework that tests an agent’s ability to handle real-world ML development scenarios. Our benchmark is particularly relevant,, as ML development increasingly relies on large language models and AI agents to assist developers. Understanding the capabilities and limitations of these agents in handling practical ML development tasks is crucial for their effective deployment in production environments.
@@ 3 Benchmark Design
ML-Dev-Bench comprises of 30 carefully designed tasks that evaluate various aspects of ML development. These tasks are structured to assess both specific technical capabilities (like handling datasets, model implementation) and broader problem-solving skills (like model training and performance improvement) that are essential in real-world ML development. The tasks span several key categories of ML development shown in Table 1:
@@
    Model Performance tasks challenge agents to improve baseline implementations through iterative experimentation and hypothesis testing.
Category ReAct-Sonnet OH-Sonnet OH-Gemini Aide-4o ReAct-4o
Dataset Handling 100% (3/3) 100% (3/3) 66% (2/3) 33% (1/3) 0% (0/3)
API Integration 100% (1/1) 100% (1/1) 0% (0/1) 0% (0/1) 100% (1/1)
Model Training 67% (4/6) 83% (5/6) 33% (2/6) 33% (2/6) 50% (3/6)
Debugging 57% (4/7) 57% (4/7) 14% (1/7) 29% (2/7) 14% (1/7)
Model Implementation 29% (2/7) 29% (2/7) 0% (0/7) 0% (0/7) 0% (0/7)
Model Performance 0% (0/6) 0% (0/6) 0% (0/6) 0% (0/6) 0% (0/6)
Overall 47% (14/30) 50% (15/30) 17% (5/30) 17% (5/30) 17% (5/30)
Table 2: Category-wise Success Rates Across AI Agents with different models
@@ 4.1 Evaluation Task
Each evaluation task consists of a task description, a set of input code and data files, and a validation logic that checks the correctness of the outputs and artifacts generated by the agent. Depending on the type of task, we implement various types of validation checks including
@@ 5 Agent Setup
    Openhands: Openhands [7] is a popular open-source coding agent with state-of-the-art performance on SWE-Bench-Full [5]. We used Openhands agent v0.21.1 and customized the runtime build to install common ML frameworks listed above. We tested the agent with Claude Sonnet 3.5 10-2022 model which is the current best performing model with the agent on SWE Bench and Gemini 2.0 Flash. The agent was allowed to run for a maximum of 50 steps.
@@ 6 Performance Comparison
Performance of the agents across different task categories, Table 2 and individual tasks Table 3 reveals a consistent pattern. Performance decreases as tasks become more open-ended and complex. The success rates are highest in well-defined categories like dataset handling and basic debugging with clear instructions, but drop significantly in open-ended and long-running tasks like model performance tasks where no agent succeeded.
OpenHands-Sonnet (OH-Sonnet) and ReAct-Sonnet are the two best performing agents performing agents with 50% and 47% success rate respectively, while OH-Gemini, AIDE-4o and React-4o achieve 17% success rate.
Across the 14 common successful tasks between OH-Sonnet and ReAct-Sonnet, we did not observe any discernible trends in cost. In long running tasks, OH-Sonnet used more tokens on average, especially in debugging and training tasks, while ReAct had higher usage in model implementation tasks like ChannelViT. Even in tasks with higher token usage, the costs dont scale proportionally due to efficient prompt caching in Openhands.
@@ 6.1 ReAct-Sonnet
ReAct-Sonnet had a success rate of 47% (14/30 tasks), demonstrating good performance in specific, well-defined tasks like dataset handling (3/3 tasks), basic model training (4/6 tasks), and debugging tasks with clear specifications (4/7 tasks). However, its performance degraded significantly in model implementation tasks where it needed to modify multiple files and in model performance tasks which are open-ended and long-running. The agent had some common failure modes we list below:
@@
    Premature task termination: In long-running training scenarios, it doesnt wait until completion of tasks but returns control saying training should complete and achieve required performance.
@@
    In long running tasks it fails to successfully implement sub-tasks which it handled correctly in isolation. For example, it correctly handles the noisy imagenette dataset setup in dedicated download tasks but fails the same operation when it’s part of the longer training pipeline.
6.2 OpenHands-Sonnet
OpenHands-Sonnet demonstrated the highest success rate at 50% (15/30 tasks), showing robust performance across most categories. The agent successfully completed all dataset handling tasks (3/3) and showed strong performance in model training (5/6) and debugging (4/7).
The agent particularly excelled in structured tasks and showed better persistence in long-running operations compared to ReAct-Sonnet. However, it still failed to complete any of the model performance tasks (0/5), indicating limitations in open-ended problem-solving scenarios requiring iterative improvement.
6.3 OpenHands-Gemini
OpenHands-Gemini achieved a 17% success rate (5/30 tasks), demonstrating limited capabilities across most evaluation categories. The agent showed some competence in basic dataset handling tasks and setting up pretrained models for training purposes. However, it performed suboptimally in debugging (1/7), model implementation (0/7), and performance improvement tasks (0/6). The agent exhibited several consistent failure modes. It frequently failed to follow task instructions regarding creation of output artifacts. In debugging scenarios, it often modified test files despite explicit instructions not to do so. The agent also demonstrated significant limitations in file editing capabilities and fixing code errors.
@@ 6.4 ReAct-4o
ReAct-4o with 17% success rate (5/30 tasks) had some success with tasks with well-defined specifications (WandB logging, downloading a specific model from Torchvision), and certain debugging tasks. However it did struggle on other tasks in the same categories. It also failed on the relatively easier tasks like dataset download due to not following instructions, ran into indentation errors while attempting to debug code and failed to produce output artifacts as required by certain tasks.
@@ 6.5 Aide-4o
Aide-4o had a 17% success rate (5/30 tasks), demonstrating limitations across most categories. The agent managed to complete some basic dataset handling and debugging tasks but struggled with model training (2/6) and completely failed in model implementation and model performance categories.
@@ 7 Conclusion
We presented ML-Dev-Bench, a benchmark focused on ML development workflows consisting of 30 tasks. We evaluated 3 agents on this benchmark - ReAct (with Claude Sonnet and GPT-4o), Openhands and AIDE; Openhands with Claude Sonnet performed the best out of these. Our open-source evaluation framework provides a foundation for the community to build upon. We encourage researchers and practitioners to explore various directions: analyzing the impact of scaling compute on these agents, studying variance in success metrics across multiple runs, evaluating emerging reasoning models like DeepSeek-R1 and o1/o3, and expanding the problem categories to include areas such as label and data collection. We welcome contributions to enhance the benchmark’s scope, methodology, and agent evaluation pipeline.